### v1

@@ 0.A.2 Implementation and Details and Experimental Settings
1:initialize DOrA’s model weights $\phi$.
@@
6:     use Algorithm A1 with $\{P,L,B\}$ as inputs to synthesize $\{D^{aug},O^{aug}_{1:B},p^{aug}_{1:B}\}$.
@@ 0.A.2.1 NR3D
For NR3D, we use pre-trained Pointnet++ to classify all object proposals as object labels $L$ following BUTD-DETR [21]. We pre-train DOrA for 15k steps on ScanNet scene point cloud and our augmented samples in Sec. 3.4 and continue on real-world pairs in NR3D for 120k steps using a batch size of 24.
0.A.2.2 ScanRefer
For ScanRefer, since ground-truth object proposals are unavailable, we adopt the visual encoder of M3DRef-CLIP [52] that adopts PointGroup [22] to perform object segmentation as proposals $P$ and object classification as labels $L$. However, We do not perform object-aware pre-training for ScanRefer since imperfect object proposals may lead to noisy synthetic samples and therefore hinder the training stability of DOrA.
0.A.2.3 Baselines
@@
For baselines [2, 17, 46, 21, 20, 5, 8, 53, 31, 52, 44] on NR3D and ScanRefer in Table 1,  2, and  3, we utilize their official public implementations with different amounts of available training samples and the full testing set to evaluate their low-resource performance. For the full-data ($100$%) scenario, we acquire their performance either on the official leaderboard of NR3D/ScanRefer or their published papers. Especially, the implementation of CORE-3DVG [44] is currently unavailable, so we only report its full-data ScanRefer performance, which is presented in their paper. On the other hand, CoT3DRef [5] does not release the parsed datasets required to train their referential-order-based framework, so we only report its performance under settings of full- and 10% training data, which are presented in their original paper.
@@
Appendix 0.B Prefix Prompt Examples of Two-Stage ICL for Deriving Referential Order
@@
This section lists the complete prompts of our two-stage ICL using GPT-3.5-Turbo [48] for referential order generation in Sec. 3.2. Also, we exhibit parsing results of 4 samples in the NR3D testing set and compare with CoT3DRef [5] that also establishes the extraction and usage of referential order.
0.B.1 Prefix Prompt of First-Stage ICL
@@
The first-stage prompt in Sec. 3.2 is used to acquire the summarized description $D^{{}^{\prime}}$ and target object $O_{B}$ of the original description. Our first-stage prompt is as follows:
@@
I have some descriptions, each describing a specific target object in a room. However, they may have some redundant clauses or words. Your task is to summarize them into a shorter description. Also tell me what is the target object.
@@
description 1: Assume you are facing the door in the room, find the larger cabinet to its left.
@@
description 3: In the bedroom, you will see a sheer curtain, beside the curtain is the steel window you need to find.
@@
description 4: Please find the towel hanging on the wall in the bathroom with other three towels. You should find the one which is nearest to the door. Or say it is on the door’s right side.
@@
summarized description 6: The bookshelf to the right of the clock facing a cabinet.
@@
description 8: The purple pillow on the right side of the bed on when facing it. Not the one the left side and the one on the middle of the bed.
@@
description 9: The brown door at the end of the living room, next to the trash cans which is full of garbage.
@@
description 10: The shoes that is placed in the middle of five shoes near the door in the room.
@@
0.B.2 Prefix Prompt of Second-Stage ICL
The second-stage prompt in Sec. 3.2 is used to acquire the referential order $O_{1:B}$ based on the target object $O_{B}$ and the summarized description $D^{\prime}$. The complete prompt is as follows:
I have some descriptions, each describing a specific target object with some supporting anchor objects helping the localization. We can find the specific target object by tracing the referential order of anchor objects step by step. Your task is to provide a correct referential order. Also tell me what are the mentioned anchor objects.
@@ 0.B.2 Prefix Prompt of Second-Stage ICL
description 3: The trash can that is on the right of the king size bed.
@@
description 10: Please find the towel hanging on the wall in the bathroom with other three towels. You should find the one nearest to the door. Or say it is on the door’s right side.
@@
Now for the description: [DESCRIPTION], give me the anchor objects and the referential order. Your answer must be in the form "referential order: , anchor objects: ".
0.B.3 Examples of Derived Referential Order
@@
To show that our two-stage ICL produces reasonable referential orders, we provide examples and comparisons between ours and CoT3DRef [5]’s parsing results. Note that CoT3DRef currently does not release their complete parsing results of NR3D, so we manually leverage its released prompt files in their paper and official repository to query the GPT-3.5-Turbo. We display 4 examples in Tab. A5, where CoT3DRef misses an anchor object “bed” in the third example and includes a redundant object “shelves” as an anchor object in the fourth example, while our DOrA produces proper results in both cases. This shows the effectiveness of two-stage ICL strategy.
@@
Table A5: Examples of LLM-parsed referential order from CoT3DRef [5] and DOrA. Note that the blue text represent the ideal anchor/target objects and the red text represent the redundant object that should not appear in the referential order. We can see that MVT misses one anchor object in the third example and includes a redundant object in the fourth example, while DOrA predicts proper order for both cases.
Description
CoT3DRef [5]
DOrA (Ours)
The pillow closest to the foot of the bed.
bed -> pillow
bed -> pillow
@@
Facing the bed, it’s the large white pillow on the right. The second one from the headboard.
bed -> headboard -> pillow
bed -> headboard -> pillow
The front pillow on the bed with the laptop.
bed -> pillow
laptop -> bed -> pillow
The window near the table, not the one near the shelves.
table -> shelves -> window
table -> window
Appendix 0.C Ablation Studies on Training Objectives
This section performs the ablation study on several training objectives used for DOrA mentioned in Sec. 3. In particular, we investigate the influence of $\mathcal{L}_{mask}$, $\mathcal{L}_{crd}$, and $\mathcal{L}_{text}$ for low-resource (1% training data) and full-resource (100% training data) scenarios, as shown in Tab. A6. Note that we only conduct target object classification on the text feature $T$ when deactivating $\mathcal{L}_{text}$, and apply classification on both anchor and target objects when adopting $\mathcal{L}_{text}$, as mentioned in Sec. 3.3. Also, for all experiments in Tab. A6, the pre-training with synthetic data in Sec. 3.4 is applied for fair comparison. In summary, $\mathcal{L}_{mask}$, $\mathcal{L}_{crd}$, and $\mathcal{L}_{text}$ all impose positive effects on models’ performances under the two settings, verifying DOrA’s design in Sec. 3.
[…]
31.1 54.2
31.6 54.8
32.4 56.0
33.5 59.7
Appendix 0.D Model Performance Under Different Order Length
This section explores DOrA’s performance gain for descriptions with different order lengths compared with MVT [20] to show the effectiveness of our design of progressive location to the target object. We split the original NR3D testing set into subsets that possess order lengths of 1, 2&3, and 4&5. An order length of 1 (1092 samples in the testing set) means that only the target object is mentioned in the description. Samples with an order length of 2 or 3 (6184 samples in the
Table A7: Grounding accuracy on NR3D subsets regarding different parsed referential order lengths. Note that referential order length $=1$ means only the target object is mentioned in the description. The improvement of DOrA over MVT [20] grows as the parsed order length increases.
Method Order Length overall
1 2&3 4&5
MVT [20] 59.4 55.0 46.7 55.1
DOrA (Ours) 61.6 (+2.2) 59.6 (+4.6) 52.0 (+5.3) 59.7 (+4.6)
testing set) have 1 or 2 anchor objects mentioned other than the target object. Similarly, samples with an order length of 4 or 5 (209 samples in the testing set) have 3 or 4 anchor objects mentioned other than the target object. A longer referential order may generally denote a longer and more complicated description. As shown in Tab. A7, though only a 2.2% performance gain is obtained for target-only samples, DOrA is more advantageous when dealing with lengthy descriptions, with 5.3% performance gain achieved for descriptions with order length of 4 or 5.
Appendix 0.E Visualization of Responses in Each Referring Block
Figure A5: Visualization of Responses in Each Referring Block. An example in NR3D is shown in this figure. We color each object proposal according to their feature response in $F_{1:(B+1)}$, where larger response with brighter color. Note that the blue bounding box represents the ground truth target object. We can see that our DOrA progressively locate the target object by considering the referential order by first focus the wall then the screen and finally the trash can.
To show that our DOrA indeed progressively locates the target object following the derived referential order, we visualize the feature response of $F_{1:(B+1)}$ ($B=4$) in Fig. A5. The blue bounding box indicates the ground-truth target object, and we color the object proposal according to the response of their corresponding features in $F_{1:(B+1)}$, where a brighter color represents a higher response. We can see that the responses to object proposals are originally cluttered. As our referential blocks are applied, the response of anchor/target objects becomes larger and finally locates the ideal target object in the last feature $F_{5}$.

### v2

@@ A.2 Implementation and Details and Experimental Settings
1:initialize Vigor’s model weights $\phi$.
@@
6:    use A1 with $\{P,L,B\}$ as inputs to synthesize $\{D^{aug},O^{aug}_{1:B},p^{aug}_{1:B}\}$.
@@ A.2.1 NR3D
For NR3D, we use pre-trained Pointnet++ to classify all object proposals as object labels $L$ following BUTD-DETR [21]. We warm-up Vigor for 15k steps on ScanNet scene point cloud and our augmented samples in Sec. 3.4 and continue on real-world data pairs in NR3D (around 1.2k, 12k, and 120k steps for 1%, 10%, and 100% data, respectively) using a batch size of 24. With one NVIDIA Tesla V100 GPU, the warm-up takes around 3 hours and the NR3D real-world data training takes around 24 hours when training on 100% data.
A.2.2 Baselines
@@
For baselines [2, 17, 46, 21, 20, 5, 8, 53, 31, 52, 44] on NR3D and ScanRefer in Table 1,  2, and  3, we utilize their official public implementations with different amounts of available training samples and the full testing set to evaluate their low-resource performance. For the full-data ($100$%) scenario, we acquire their performance either on the official leaderboard of NR3D/ScanRefer or their published papers. Especially, the implementation of CORE-3DVG [44] is currently unavailable, so we only report its full-data ScanRefer performance, which is presented in their paper.
Appendix B Additional Quantitative Results
B.1 More Results on NR3D
Figure A5: Quantitative results on NR3D We can see that when the amount of data is relatively many (above 30%), our Vigor is comparable to MVT+CoT3DRef [5]. However, as the amount of data reduces, our Vigor performs better over MVT-CoT3DRef.
In Sec. 4, we have shown comparisons between Vigor and several state-of-the-arts under data efficient scenarios (1% $\sim$ 10% of data). Here, we provide a comparison between Vigor and MVT+CoT3DRef [5] from 1% to 100% of data in Figure A5. In this figure, Vigor outperforms CoT3DRef with a noticeable margin when the amount of data is limited and is comparable to CoT3DRef when the data amount is over 30%, showing that our Vigor is suitable under different settings.
B.2 Experiments on ScanRefer
B.2.1 Dataset and setting
ScanRefer [8] contains 36665/9508 description-target pairs across a total of 800 indoor scenes in its training/validation set, where the descriptions are also collected from human annotators. Also derived from ScanNet [12] but different from NR3D, perfect object proposals are not available in ScanRefer, and therefore, additional object proposers are required for all methods. Nevertheless, Acc in % under 0.25 and 0.5 intersection over union (IoU) are used as the metrics for ScanRefer. In ScanRefer, since ground-truth object proposals are unavailable, we adopt the visual encoder of M3DRef-CLIP [52] that adopts PointGroup [22] to perform object segmentation as proposals $P$ and object classification as labels $L$. However, we do not perform object-aware pre-training for ScanRefer since imperfect object proposals may lead to noisy synthetic samples and hinder Vigor’s training stability.
B.2.2 Quantitative Results
The results are shown in Table A4. We can see that our Vigor demonstrates slightly preferable results with only 5% and 10% of data compared to other approaches. It is comparable to SOTA methods when given 100% data. Note that we only report CORE-3DVG’s [44] results on 100% (36665) ScanRefer training samples since its implementation is currently unavailable.
Table A4: Grounding accuracy (%) on the ScanRefer validation set. In this table, different amounts of training data are considered. Note that the official implementation of CORE-3DVG [44] has not yet been released; we show the numbers in their original paper in this table.
Method Labeled Training Data
5% 10% 100%
Acc@0.25 Acc@0.5 Acc@0.25 Acc@0.5 Acc@0.25 Acc@0.5
ScanRefer [8] 23.0 12.0 27.4 15.0 35.5 22.4
3DVG-Trans. [53] 35.3 23.3 39.0 29.0 45.9 34.5
3D-SPS [31] 28.4 16.9 32.9 22.9 48.8 37.0
BUTD-DETR [21] 38.2 26.2 40.3 28.3 52.2 39.8
M3DRef-CLIP [52] 37.2 29.9 40.0 32.4 52.5 44.7
CORE-3DVG [44] - - - - 56.8 43.8
Vigor (Ours) 39.8 31.5 43.6 34.9 52.8 44.8
Appendix C Ablation Studies on Training Objectives
This section performs the ablation study on several training objectives used for Vigor mentioned in Sec. 3. In particular, we investigate the influence of $\mathcal{L}_{mask}$, $\mathcal{L}_{crd}$, and $\mathcal{L}_{text}$ for low-resource (1% training data) and full-resource (100% training data) scenarios, as shown in A5. Note that we only conduct target object classification on the text feature $T$ when deactivating $\mathcal{L}_{text}$, and apply classification on both anchor and target objects when adopting $\mathcal{L}_{text}$, as mentioned in Sec. 3.3. Also, for all experiments in A5, the pre-training with synthetic data in Sec. 3.4 is applied for fair comparison. In summary, $\mathcal{L}_{mask}$, $\mathcal{L}_{crd}$, and $\mathcal{L}_{text}$ all impose positive effects on models’ performances under the two settings, verifying Vigor’s design in Sec. 3.
[…]
31.1 54.2
31.6 54.8
32.4 56.0
33.5 59.7
Appendix D Model Performance Under Different Order Length
(a) Grounding accuracy with different maximum (padded) referential order length $B$. Note that $B$ is equal to and denoted as the number of Object-Referring blocks.
Figure A6: (a) Grounding accuracies with varying Object-Referring block numbers $B$ and (b) statistics of referential order length of NR3D. It can be seen that the model performance saturates at $B$=4, matching length statistics of NR3D.
In this section, we analyze the effectiveness of padded order length $B$ of $O_{1:B}$ (and also the number of Referring Blocks) in Fig. A6, which illustrates the performances with different $B$ using 1% and 100% training data in Fig. 6(a). We also show the statistics of the original order length generated by our two-stage referential ordering in Fig. 6(b). We observe that although performances gradually increase with larger $B$ (i.e., considering more potential anchor objects appeared in $D$ achieves better accuracy), the performance gain saturates at $B=4$. This can be explained by looking at Fig. 6(b), which shows that only a very small amount of data exceeds an order length of $4$ in both training and testing pairs. This suggests that our selection of $B=4$ is reasonable, optimally balancing computational efficiency with prediction accuracy.
We additionally explore Vigor’s performance gain for descriptions with different order lengths compared with MVT [20] to show the effectiveness of our design of progressive location to the target object. We split the original NR3D testing set into subsets that possess order lengths of 1, 2&3, and 4&5. An order length of 1 (1092 samples in the testing set) means that only the target object is mentioned in the description. Samples with an order length of 2 or 3 (6184 samples in the
Table A6: Grounding accuracy on NR3D subsets regarding different parsed referential order lengths. Note that referential order length $=1$ means only the target object is mentioned in the description. The improvement of Vigor over MVT [20] grows as the parsed order length increases.
Method Order Length overall
1 2&3 4&5
MVT [20] 59.4 55.0 46.7 55.1
Vigor (Ours) 61.6 (+2.2) 59.6 (+4.6) 52.0 (+5.3) 59.7 (+4.6)
testing set) have 1 or 2 anchor objects mentioned other than the target object. Similarly, samples with an order length of 4 or 5 (209 samples in the testing set) have 3 or 4 anchor objects mentioned other than the target object. A longer referential order may generally denote a longer and more complicated description. As shown in A6, though only a 2.2% performance gain is obtained for target-only samples, Vigor is more advantageous when dealing with lengthy descriptions, with 5.3% performance gain achieved for descriptions with order length of 4 or 5.
Appendix E Visualization of Responses in Each Referring Block
Figure A7: Visualization of Responses in Each Referring Block. An example in NR3D is shown in this figure. We color each object proposal according to their feature response in $F_{1:(B+1)}$, where a larger response with a brighter color. Note that the blue bounding box represents the ground truth target object. We can see that our Vigor progressively locates the target object by considering the referential order by first focusing on the wall then the screen and finally the trash can.
To show that our Vigor indeed progressively locates the target object following the derived referential order, we visualize the feature response of $F_{1:(B+1)}$ ($B=4$) in A7. The blue bounding box indicates the ground-truth target object, and we color the object proposal according to the response of their corresponding features in $F_{1:(B+1)}$, where a brighter color represents a higher response. We can see that the responses to object proposals are originally cluttered. As our referential blocks are applied, the response of anchor/target objects becomes larger and finally locates the ideal target object in the last feature $F_{5}$.
@@ Appendix E Visualization of Responses in Each Referring Block
Appendix F Details and Prefix Prompt Examples of Two-Stage ICL for Deriving Referential Order
Figure A8: Referential order generation via LLM. A two-stage ICL is deployed to remove redundant information from $D$, forming a descriptive order $O_{1:B}$ for locating the target object $O_{B}$.
To have Object-Referring blocks $\{R_{1},\cdots,R_{B}\}$ in Sec. 3 to locate the target object properly, it is desirable to extract a proper referential order $O_{1:B}$ from the input description $D$. With such a referring path constructed, visual features of the associated objects can be updated for grounding purposes. This is inspired by the idea of Chain-of-Thoughts [41, 26] in LLM, as noted in CoT3DRef [5]. To achieve such a parsing task, we apply GPT-3.5-Turbo [48] as the description parser using in-context learning (ICL) [14], as depicted in Fig. A8.
However, it is not trivial to have LLM output a referential order from $D$ due to lengthy and noisy descriptions. For example, for an input description “Look at the king-size bed in the room next to a green chair, Find the pillow on the bed. Not the pillow on the sofa beside the chair.”, one would expect first to find the green chair then the king-size bed next to the chair, and finally, the pillow on the bed. Therefore, the ideal referential order is {“chair”, “bed”, “pillow”}. In the above example, the sentence “Not the pillow on the sofa which is also beside the chair.” is redundant since one can find the correct target object without this information. If we apply the LLM directly to the original description, the model may be misled by this redundant information and generate a referential order containing the object “sofa”.
To tackle the above problem, we conduct a two-stage in-context learning (ICL) scheme to remove redundant information in $D$ before producing $O_{1:B}$, as depicted in Fig. A8.
With the given target object $O_{B}$, we predict a summarized description $D^{{}^{\prime}}$ to remove redundant information in $D$ in the first stage. Then, the entire referential order $O_{1:B}$ given $D^{{}^{\prime}}$ and $O_{B}$ is produced in the second stage. For each stage of our ICL, 10 examples are provided as demonstrations of the input prompts for LLM to predict $O_{1:B}$. Due to page limitations, demonstration examples and the complete prefix prompts are presented in the supplementary materials. With $O_{1:B}$ obtained, the order-aware object referring process can be processed accordingly, as we detail next.
@@ Appendix F Details and Prefix Prompt Examples of Two-Stage ICL for Deriving Referential Order
We now list the complete prompts of our two-stage ICL using GPT-3.5-Turbo [48] for referential order generation. Also, we exhibit parsing results of 4 samples in the NR3D testing set and compare them with CoT3DRef [5] that also establishes the extraction and usage of referential order.
F.1 Prefix Prompt of First-Stage ICL
@@
The first-stage prompt is used to acquire the summarized description $D^{{}^{\prime}}$ and target object $O_{B}$ of the original description. Our first-stage prompt is as follows:
@@
I have some descriptions, each describing a specific target object in a room. However, they may have some redundant clauses or words. Your task is to summarize them into a shorter description. Also, tell me what the target object.
@@
description 1: Assume you are facing the door in the room. Find the larger cabinet to its left.
@@
description 3: In the bedroom, you will see a sheer curtain. Beside the curtain is the steel window you need to find.
@@
description 4: Please find the towel hanging on the wall in the bathroom with the other three towels. You should find the one nearest to the door. Or say it is on the door’s right side.
@@
summarized description 6: The bookshelf to the right of the clock faces a cabinet.
@@
description 8: The purple pillow on the right side of the bed when facing it. Not the one on the left side and the one in the middle of the bed.
@@
description 9: The brown door at the end of the living room, next to the trash cans, which are full of garbage.
@@
description 10: The shoes that are placed in the middle of five shoes near the door in the room.
@@
F.2 Prefix Prompt of Second-Stage ICL
The second-stage prompt is used to acquire the referential order $O_{1:B}$ based on the target object $O_{B}$ and the summarized description $D^{\prime}$. The complete prompt is as follows:
I have some descriptions, each describing a specific target object with some supporting anchor objects helping the localization. We can find the specific target object by tracing the referential order of anchor objects step by step. Your task is to provide a correct referential order. Also, tell me what the mentioned anchor objects.
@@ F.2 Prefix Prompt of Second-Stage ICL
description 3: The trash can that is on the right of the king-size bed.
@@
description 10: Please find the towel hanging on the wall in the bathroom with the other three towels. You should find the one nearest to the door. Or say it is on the door’s right side.
@@
Now for the description: [DESCRIPTION], give me the anchor objects and the referential order. Your answer must be in the form "referential order, anchor objects:. "
F.3 Examples of Derived Referential Order
@@
To show that our two-stage ICL produces reasonable referential orders, we provide examples and comparisons between ours and CoT3DRef [5]’s parsing results. In particular, we leverage CoT3DRef’s released prompt to query the GPT-3.5-Turbo. We display 4 examples in A7, where CoT3DRef misses an anchor object “bed” in the third example and includes a redundant object “shelves” as an anchor object in the fourth example, while our Vigor produces proper results in both cases. This shows the effectiveness of the two-stage ICL strategy.
@@
Table A7: Examples of LLM-parsed referential order from CoT3DRef [5] and Vigor. Note that the blue text represents the ideal anchor/target objects, and the red text represents the redundant object that should not appear in the referential order. We can see that MVT misses one anchor object in the third example and includes a redundant object in the fourth example, while Vigor predicts proper order for both cases.
Description CoT3DRef [5] Vigor (Ours)
The pillow closest to the foot of the bed. bed -> pillow bed -> pillow
@@
Facing the bed, it’s the large white pillow on the right. The second one from the headboard. bed -> headboard -> pillow bed -> headboard -> pillow
The front pillow on the bed with the laptop. bed -> pillow laptop -> bed -> pillow
The window near the table, not the one near the shelves. table -> shelves -> window table -> window
Appendix G Limitations and Social Impact
G.1 Limitations
G.1.1 LLM-parsed Referential Order
Since Vigor utilizes LLMs to generate the referential order from the input description, despite of our introduced pre-training strategy to warm the training process, the correctness of the extracted referential order would affect the training and the performance of Vigor.
G.1.2 Order Length Decision
As detailed in Sec. A, we set the length $B$ of referential order (as well as the number of referring blocks) as $4$ to conduct batch-wise training. Although we have shown that our choice of $B$ is reasonable for balancing computational efficiency and prediction accuracy in Sec. D by conducting proper experiments on NR3D, this choice appears to be dataset-specific. We leave this as a future direction to develop a more flexible architecture to be able to dynamically adjust $B$ according to the LLM-parsed referential order for each sample during training and testing.
G.2 Broader Impact
G.2.1 Applications of 3D visual grounding
Although the experiments in this paper are conducted on indoor datasets only, the task of 3D visual grounding is not restricted to indoor scenes. For the autonomous driving industry, 3D visual grounding could also be an important topic to study along with 3D object detection. As our approach is designed to promote the data efficiency of 3D visual grounding tasks, we look forward to seeing future works that consider the concept of our Vigor and apply it to autonomous driving systems since it is hard to collect enormous amounts of data for autonomous driving.
G.2.2 Potential Negative Impacts
Although the experiments in this paper show that our Vigor outperforms current SOTAs in data-efficient scenarios, one must make sure that Vigor is well-validated before applying it to a new data domain. Without properly transferring Vigor’s grounding ability to the corresponding domain, the performance could be non-ideal, introducing potential safety risks, especially in critical applications like autonomous driving.